# Compression Represents Intelligence Linearly

### v1

@@ Abstract
There is a belief that learning to compress well will lead to intelligence (Hutter, 2006). Recently, language modeling has been shown to be equivalent to compression, which offers a compelling rationale for the success of large language models (LLMs): the development of more advanced language models is essentially enhancing compression which facilitates intelligence. Despite such appealing discussions, little empirical evidence is present for the interplay between compression and intelligence. In this work, we examine their relationship in the context of LLMs, treating LLMs as data compressors. Given the abstract concept of “intelligence”, we adopt the average downstream benchmark scores as a surrogate, specifically targeting intelligence related to knowledge and commonsense, coding, and mathematical reasoning. Across 12 benchmarks, our study brings together 30 public LLMs that originate from diverse organizations. Remarkably, we find that LLMs’ intelligence – reflected by average benchmark scores – almost linearly correlates with their ability to compress external text corpora. These results provide concrete evidence supporting the belief that superior compression indicates greater intelligence. Furthermore, our findings suggest that compression efficiency, as an unsupervised metric derived from raw text corpora, serves as a reliable evaluation measure that is linearly associated with the model capabilities. We open-source our compression datasets as well as our data collection pipelines to facilitate future researchers to assess compression properly. [^1] https://github.com/hkust-nlp/llm-compression-intelligence.
@@ 1 Introduction
We measure intelligence along three key abilities: knowledge and commonsense, coding, and mathematical reasoning. Specifically, we begin by collecting external raw corpora in the respective domain (e.g. GitHub code for the coding ability) and evaluate the efficiency of various LLMs in compressing the corpus. Subsequently, we assess these models on a range of downstream tasks, using the average benchmark scores to gauge their domain-specific intelligence. The goal is to examine the correlation between the models’ downstream ability and their compression efficiency. On experiments across 30 public LLMs and 12 diverse benchmarks, we uncover a remarkable finding: LLMs’ downstream ability is almost linearly correlated with their compression efficiency, with a Pearson correlation coefficient of around -0.95 for each evaluated domain of intelligence as shown in Figure 1. Notably, this linear relationship even extends to most individual benchmarks as well.
Recent and concurrent works have explored the relationship between benchmark scores and compression-equivalent metrics like validation loss within the same model series, where the model checkpoints share most configurations such as model designs, tokenizers, and data (Wei et al., 2023; Gadre et al., 2024; Du et al., 2024). Our study, however, is the first to document a linear correlation between compression and intelligence in LLMs across varying model sizes, tokenizers, context window lengths, and pretraining data distributions. Our findings establish the linear correlation between compression and intelligence as a universal principle, providing empirical support for the longstanding belief that superior compression is indicative of greater intelligence.
@@ 3.3 Overfitting Benchmarks
LLMs may be overoptimized towards certain benchmarks, in this case, the benchmark scores are not a good proxy for intelligence anymore.333“When a measure becomes a target, it ceases to be a good measure” – Goodhart’s law. For example, Bi et al. (2024) show that when adding multi-choice QA training data, the model is able to achieve over 10 absolute points improvement on multi-choice QA benchmarks such as MMLU, while its performance on general QA benchmarks like TriviaQA remains unchanged. In the math domain, recent researches find that some models heavily optimize towards the GSM8K and MATH benchmarks, while performing poorly on held-out math tests (Paster, 2023). In such scenarios, these benchmark scores are not reliable enough to represent the models’ intelligence in the respective area. Although we focus on base models in this work, where such overfitting phenomenon may be less common, it is still possible that some LLMs are specially trained on the training data of the corresponding benchmarks in the pretraining stage, or even worse, suffer from test data contamination issues as evidenced in Wei et al. (2023). To identify such cases, we adopt the MIN-KPROB method (Shi et al., 2024) which is proposed to detect whether the given text is exposed to the model during pretraining. The MIN-KPROB approach selects the $k\%$ tokens in a given example with minimum probabilities, if the average probability of these tokens is high, then the example is likely to be present in the pretraining data. We define the MIN-KScore to be the average negative log-likelihood of these selected tokens. In the experiments next, we compute the MIN-KScore for all the test splits of the benchmarks, and for the training splits as well if available. We will spot and discuss the LLMs with extremely abnormal MIN-KScores. We choose $k$ to be 20 in our experiments as suggested by Shi et al. (2024).
@@ 4 How is Intelligence Correlated with Compression?
To provide a comprehensive study of the correlation between compression efficiency and intelligence, we include 8 series of general-purpose language models, covering diverse organizations and varying in size, as shown in Table 2. Notably, in addition to dense transformer models, we further include Mixtral-8x7B (MistralAI, 2023), which utilizes a sparse mixture-of-experts (MoE) architecture. Furthermore, we incorporate state-of-the-art models specialized in coding and mathematical reasoning in addition to general-purpose models, as outlined in Table 2. We assess general-purpose LLMs across all benchmarks, while we include code LLMs in the code area only and mathematical LLMs in the mathematical reasoning area only. We note that our investigation focuses exclusively on well-trained LLMs, implying that the intelligence evaluated is likely already manifest, in contrast to models where such abilities have yet to emerge (Wei et al., 2022). This is because LLMs that are not optimized enough tend to perform poorly, where the results are overly noisy and less meaningful, complicating the analysis. Models are evaluated using few-shot in-context learning or in a zero-shot manner, adhering to the established norms for the respective benchmarks. All the models are evaluated exactly the same way with the same codebase for a fair comparison. More details on evaluation are included in Appendix A.3.
@@ 4.1 Main Results – Compression Represents Intelligence Linearly
We have summarized our main results in Figure 1, where we examine the overall relationship by computing the average benchmark scores across all three areas, and the average BPC across all the compression corpora. The average score and the average BPC demonstrate a highly linear correlation qualitatively in the visualization. Quantitative results with a Pearson correlation coefficient of -0.94 and an RMSE of 2.8% further verify the linear relationship. We consider such a linear correlation significant given the noisy nature of evaluating on downstream benchmarks – it is known that varying the prompts or decoding hyperparameters could easily cause several points of difference for the task scores (Lu et al., 2022; Gonen et al., 2022; Sclar et al., 2024). We also report the results for the three key areas respectively which demonstrate similar phenomena qualitatively and quantitatively. We note that several previous and concurrent works have studied the correlation between benchmark scores and validation losses in limited settings – they focus on either the model checkpoints of the same model over the course of pretraining or the same model series (Lu et al., 2023; Wei et al., 2023; Gadre et al., 2024; Du et al., 2024). Their studied checkpoints share designs and tokenizers and are pretrained on the same data distributions. However, our work escapes the constraints on the same model (series) and compares across models that diverge dramatically on tokenizers, model designs, and pretraining data. Our results mark the linear correlation between compression and intelligence in a general manner, establishing it as universal principle. The specific benchmark scores and BPC for every model are reported in Appendix B. Next, we discuss the results in the three areas with more details respectively, and extend the examination to individual benchmarks as well.
@@
We present both the average and individual benchmark results for the knowledge and commonsense area in Figure 2. Notably, the scores from individual benchmarks exhibit a strong linear correlation with compression efficiency, indicated by Pearson coefficients of -0.877, -0.947, and -0.888 for HellaSwag, TriviaQA, and MMLU, respectively. The Root Mean Square Error $e$ is approximately 2 absolute points, which falls within the expected range of variation – Minor fluctuations in evaluation details, such as prompt adjustments, can readily account for a 2-point accuracy shift (Sclar et al., 2024). Meanwhile, we observe that the linear correlation in both HellaSwag and MMLU is not as pronounced as in TriviaQA, albeit for different reasons. For HellaSwag, the average accuracy among models has exceeded 0.8, hinting at a saturation point in performance that blurs the distinctions between models. Conversely, MMLU, which draws heavily from real-world examinations and textbooks, presents a slight mismatch in distribution compared to the Common Crawl dataset used to measure compression efficiency. Further experiments on MMLU and textbook compression are detailed in Appendix E, where a Pearson coefficient of -0.926 is observed. We note that for MMLU evaluation, we employ a cloze-style input format for this popular multiple-choice question (MCQ) benchmark, presenting only the question without the options, and subsequently selecting the choice with the highest likelihood. We intentionally opted for this less conventional evaluation style to minimize the risk of models overfitting to the MCQ format, thus over-optimizing the scores. Please see Appendix D for detailed discussion and ablation.
Coding:
Similar to the knowledge and commonsense area, there is a strong linear correlation between the models’ compression efficiency and its coding ability, with the Pearson correlation coefficient within $[-1,-0.9]$ for each individual benchmark. Furthermore, despite being of different architectures, the MoE model, Mixtral-8x7B, still fits well into this linear correlation. We notice the strongest linear correlation within the DS-1000 benchmark, with a Pearson coefficient of -0.958, surpassing those of HumanEval and MBPP. We think that this is because we collect Python code from the popular GitHub repositories in recent months, which are mostly related to data science and machine learning, aligned closer with DS-1000. We note that the Deepseek-coder model series are consistently well-above the linear fit, demonstrating strong abilities on the HumanEval and MBPP benchmarks. We hypothesize that this is due the fact that Deepseek-coder models are exposed to private, task-specific data during pretraining that help HumanEval and MBPP significantly, yet we do not have reliable methods to detect the existance of such data and verify this hypotheis.
@@ 4.2 Remaining Questions
As an example, we investigate the effect of size of compression corpus on the knowledge and commonsense ability. We vary the number of characters of the compression corpus from 50K to 100M, observing changes in the Pearson correlation coefficient and the RMSE depicted in Figure 7. As the number of sampled characters increases, the linear correlation gradually strengthens and reaches saturation at around 30M. This observation suggests that the compression corpus is sufficiently large with only tens of millions of characters. We further provide an analysis of the impact of random sampling on compression evaluation in Appendix C.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ A.3 Benchmark Evaluation
MATH 4 Generation-based AGIEval Acc
@@
Table 3: Evaluation details of benchmarks. The perplexity-based evaluation method involves calculating and normalizing the perplexity for each option and selecting the option with the lowest perplexity. For the majority of benchmarks, we utilize the LM-Evaluation-Harness framework (Gao et al., 2023a). Additionally, we incorporate other sources as supplementary, including Deepseek-Coder (Guo et al., 2024), InCoder (Fried et al., 2023) and AGIEval (Zhong et al., 2023).
@@ 6-7B Parameter Model
Deepseek-llm-7b 17.2 4.0 31.8 37.8 22.7 0.500
Deepseek-math-7b 59.2 21.5 60.0 50.5 47.8 0.420
Falcon-7b 3.9 1.8 29.5 34.7 17.5 0.541
Llama-1-7b 10.7 2.6 26.4 36.1 18.9 0.510
Llama-2-7b 13.7 3.1 28.6 39.5 21.2 0.500
Llemma-7b 31.7 10.8 52.7 39.6 33.7 0.431
Mistral-7b 33.4 9.2 48.6 46.1 34.3 0.443
Qwen-7b 42.8 10.4 43.6 41.3 34.5 0.483
Yi-6B 31.1 4.9 33.6 40.8 27.6 0.483
@@
Llama-1-13b 13.6 3.6 28.6 39.0 21.2 0.487
Llama-2-13b 21.9 4.2 31.4 42.8 25.1 0.475
Qwen-14b 55.7 19.2 57.3 47.1 44.8 0.450
@@
Falcon-40b 23.0 3.6 31.8 43.0 25.4 0.482
Llama-1-30b 32.1 5.8 36.4 44.9 29.8 0.456
Llemma-34b 46.2 15.6 54.5 48.0 41.1 0.416
Yi-34B 61.4 13.0 52.3 50.8 44.4 0.421
@@
Deepseek-llm-67b 56.0 10.7 54.1 50.1 42.7 0.430
Llama-1-65b 42.5 8.0 41.8 45.3 34.4 0.441
Llama-2-70b 50.4 9.9 50.9 48.0 39.8 0.429
Mixtral-8x7b 53.4 17.3 64.1 54.4 47.3 0.394
Qwen-72b 72.6 25.9 71.4 52.3 55.5 0.415
@@ Appendix E Better Alignment between MMLU and Textbooks
The MMLU benchmark offers a comprehensive evaluation across various domains and tasks, leveraging datasets sourced from real-world examinations and textbooks. As a result, textbooks from various educational levels closely align with the benchmark’s distribution. We utilize a small collection of textbooks sourced from the internet as validation data. The results are shown in Figure  11. A stronger linear correlation is observed between the compression performance on the textbook data and the MMLU accuracy, characterized by a high Pearson correlation coefficient of -0.926.

### v2

@@ Abstract
There is a belief that learning to compress well will lead to intelligence (Hutter, 2006). Recently, language modeling has been shown to be equivalent to compression, which offers a compelling rationale for the success of large language models (LLMs): the development of more advanced language models is essentially enhancing compression which facilitates intelligence. Despite such appealing discussions, little empirical evidence is present for the interplay between compression and intelligence. In this work, we examine their relationship in the context of LLMs, treating LLMs as data compressors. Given the abstract concept of “intelligence”, we adopt the average downstream benchmark scores as a surrogate, specifically targeting intelligence related to knowledge and commonsense, coding, and mathematical reasoning. Across 12 benchmarks, our study brings together 31 public LLMs that originate from diverse organizations. Remarkably, we find that LLMs’ intelligence – reflected by average benchmark scores – almost linearly correlates with their ability to compress external text corpora. These results provide concrete evidence supporting the belief that superior compression indicates greater intelligence. Furthermore, our findings suggest that compression efficiency, as an unsupervised metric derived from raw text corpora, serves as a reliable evaluation measure that is linearly associated with the model capabilities. We open-source our compression datasets as well as our data collection pipelines to facilitate future researchers to assess compression properly. [^1] https://github.com/hkust-nlp/llm-compression-intelligence.
@@ 1 Introduction
We measure intelligence along three key abilities: knowledge and commonsense, coding, and mathematical reasoning. Specifically, we begin by collecting external raw corpora in the respective domain (e.g. GitHub code for the coding ability) and evaluate the efficiency of various LLMs in compressing the corpus. Subsequently, we assess these models on a range of downstream tasks, using the average benchmark scores to gauge their domain-specific intelligence. The goal is to examine the correlation between the models’ downstream ability and their compression efficiency. On experiments across 31 public LLMs and 12 diverse benchmarks, we uncover a remarkable finding: LLMs’ downstream ability is almost linearly correlated with their compression efficiency, with a Pearson correlation coefficient of around -0.95 for each evaluated domain of intelligence as shown in Figure 1. Notably, this linear relationship even extends to most individual benchmarks as well.
Recent and concurrent works have explored the relationship between benchmark scores and compression-equivalent metrics like validation loss within the same model series, where the model checkpoints share most configurations such as model designs, tokenizers, and data (Xia et al., 2023; Wei et al., 2023; Gadre et al., 2024; Du et al., 2024). Our study, however, is the first to document a linear correlation between compression and intelligence in LLMs across varying model sizes, tokenizers, context window lengths, and pretraining data distributions. Our findings establish the linear correlation between compression and intelligence as a universal principle, providing empirical support for the longstanding belief that superior compression is indicative of greater intelligence.
@@ 3.3 Overfitting Benchmarks
LLMs may be overoptimized towards certain benchmarks, in this case, the benchmark scores are not a good proxy for intelligence anymore.333“When a measure becomes a target, it ceases to be a good measure” – Goodhart’s law. For example, Bi et al. (2024) show that when adding multi-choice QA training data, the model is able to achieve over 10 absolute points improvement on multi-choice QA benchmarks such as MMLU, while its performance on general QA benchmarks like TriviaQA remains unchanged. In the math domain, recent researches find that some models heavily optimize towards the GSM8K and MATH benchmarks, while performing poorly on held-out math tests (Paster, 2023). In such scenarios, these benchmark scores are not reliable enough to represent the models’ intelligence in the respective area. Although we focus on base models in this work, where such overfitting phenomenon may be less common, it is still possible that some LLMs are specially trained on the training data of the corresponding benchmarks in the pretraining stage, or even worse, suffer from test data contamination issues as evidenced in Wei et al. (2023). To identify such cases, we adopt the MIN-KPROB method (Shi et al., 2024) which is proposed to detect whether the given text is exposed to the model during pretraining. The MIN-KPROB approach selects the $k\%$ tokens in a given example with minimum probabilities, if the average probability of these tokens is high, then the example is likely to be present in the pretraining data. In this approach, the MIN-KScore is defined to be the average negative log-likelihood of these selected tokens. In the experiments next, we compute the MIN-KScore for all the test splits of the benchmarks, and for the training splits as well if available. We will spot and discuss the LLMs with extremely abnormal MIN-KScores. We choose $k$ to be 20 in our experiments as suggested by Shi et al. (2024).
@@ 4 How is Intelligence Correlated with Compression?
To provide a comprehensive study of the correlation between compression efficiency and intelligence, we include 9 series of general-purpose language models, covering diverse organizations, varying in size and architectures, as shown in Table 2. Notably, in addition to dense transformer models, we further include Mixtral-8x7B (MistralAI, 2023) and Jamba-52B (Lieber et al., 2024) to validate our findings across different architectures. Mixtral-8x7B utilizes a sparse mixture-of-experts (MoE) architecture, whereas Jamba-52B is based on a hybrid Transformer-Mamba MoE architecture. Furthermore, we incorporate state-of-the-art models specialized in coding and mathematical reasoning in addition to general-purpose models, as outlined in Table 2. We assess general-purpose LLMs across all benchmarks, while we include code LLMs in the code area only and mathematical LLMs in the mathematical reasoning area only. We note that our investigation focuses exclusively on well-trained LLMs, implying that the intelligence evaluated is likely already manifest, in contrast to models where such abilities have yet to emerge (Wei et al., 2022). This is because LLMs that are not optimized enough tend to perform poorly, where the results are overly noisy and less meaningful, complicating the analysis. Models are evaluated using few-shot in-context learning or in a zero-shot manner, adhering to the established norms for the respective benchmarks. All the models are evaluated exactly the same way with the same codebase for a fair comparison. More details on evaluation are included in Appendix A.3.
@@ 4.1 Main Results – Compression Represents Intelligence Linearly
We have summarized our main results in Figure 1, where we examine the overall relationship by computing the average benchmark scores across all three areas, and the average BPC across all the compression corpora. The average score and the average BPC demonstrate a highly linear correlation qualitatively in the visualization. Quantitative results with a Pearson correlation coefficient of -0.93 and an RMSE of 3.1% further verify the linear relationship. We consider such a linear correlation significant given the noisy nature of evaluating on downstream benchmarks – it is known that varying the prompts or decoding hyperparameters could easily cause several points of difference for the task scores (Lu et al., 2022; Gonen et al., 2022; Sclar et al., 2024). We also report the results for the three key areas respectively which demonstrate similar phenomena qualitatively and quantitatively. We note that several previous and concurrent works have studied the correlation between benchmark scores and validation losses in limited settings – they focus on either the model checkpoints of the same model over the course of pretraining or the same model series (Xia et al., 2023; Lu et al., 2023; Wei et al., 2023; Gadre et al., 2024; Du et al., 2024). Their studied checkpoints share designs and tokenizers and are pretrained on the same data distributions. However, our work escapes the constraints on the same model (series) and compares across models that diverge dramatically on tokenizers, model designs, and pretraining data. Our results mark the linear correlation between compression and intelligence in a general manner, establishing it as a universal principle. The specific benchmark scores and BPC for every model are reported in Appendix B. Next, we discuss the results in the three areas with more details respectively, and extend the examination to individual benchmarks as well.
@@
We present both the average and individual benchmark results for the knowledge and commonsense area in Figure 2. Notably, the scores from individual benchmarks exhibit a strong linear correlation with compression efficiency, indicated by Pearson coefficients of -0.881, -0.951, and -0.890 for HellaSwag, TriviaQA, and MMLU, respectively. The Root Mean Square Error $e$ is approximately 2 absolute points, which falls within the expected range of variation – Minor fluctuations in evaluation details, such as prompt adjustments, can readily account for a 2-point accuracy shift (Sclar et al., 2024). Meanwhile, we observe that the linear correlation in both HellaSwag and MMLU is not as pronounced as in TriviaQA, albeit for different reasons. For HellaSwag, the average accuracy among models has exceeded 0.8, hinting at a saturation point in performance that blurs the distinctions between models. Conversely, MMLU, which draws heavily from real-world examinations and textbooks, presents a slight mismatch in distribution compared to the Common Crawl dataset used to measure compression efficiency. Further experiments on MMLU and textbook compression are detailed in Appendix E, where a Pearson coefficient of -0.930 is observed. We note that for MMLU evaluation, we employ a cloze-style input format for this popular multiple-choice question (MCQ) benchmark, presenting only the question without the options, and subsequently selecting the choice with the highest likelihood. We intentionally opted for this less conventional evaluation style to minimize the risk of models overfitting to the MCQ format, thus over-optimizing the scores. Please see Appendix D for detailed discussion and ablation.
@@
Coding:
Similar to the knowledge and commonsense area, there is a strong linear correlation between the models’ compression efficiency and its coding ability, with the Pearson correlation coefficient within $[-1,-0.9]$ for each individual benchmark. Furthermore, despite having significantly different architectures, both Mixtral-8x7B and Jamba-52B fit well into this linear correlation, demonstrating that our findings are agnostic to model architecture. We notice the strongest linear correlation within the DS-1000 benchmark, with a Pearson coefficient of -0.946, surpassing those of HumanEval and MBPP. We think that this is because we collect Python code from the popular GitHub repositories in recent months, which are mostly related to data science and machine learning, aligned closer with DS-1000. We note that the Deepseek-coder model series are consistently well-above the linear fit, demonstrating strong abilities on the HumanEval and MBPP benchmarks. We hypothesize that this is due the fact that Deepseek-coder models are exposed to private, task-specific data during pretraining that help HumanEval and MBPP significantly, yet we do not have reliable methods to detect the existance of such data and verify this hypotheis.
@@ 4.2 Remaining Questions
As an example, we investigate the effect of size of compression corpus on the knowledge and commonsense ability. We vary the number of characters of the compression corpus from 50K to 100M, observing changes in the Pearson correlation coefficient and the RMSE depicted in Figure 7. As the number of sampled characters increases, the linear correlation gradually strengthens and reaches saturation at around 3M. This observation suggests that the compression corpus is sufficiently large with only tens of millions of characters. We further provide an analysis of the impact of random sampling on compression evaluation in Appendix C.
@@ A.3 Benchmark Evaluation
MATH 4 Generation-based OpenCompass Acc
@@
Table 3: Evaluation details of benchmarks. The perplexity-based evaluation method involves calculating and normalizing the perplexity for each option and selecting the option with the lowest perplexity. For the majority of benchmarks, we utilize the LM-Evaluation-Harness framework (Gao et al., 2023a). Additionally, we incorporate other sources as supplementary, including Deepseek-Coder (Guo et al., 2024), InCoder (Fried et al., 2023), OpenCompass (Contributors, 2023) and AGIEval (Zhong et al., 2023).
@@ 6-7B Parameter Model
Deepseek-llm-7b 17.2 4.5 31.8 37.8 22.8 0.500
Deepseek-math-7b 59.2 28.0 60.0 50.5 49.4 0.420
Falcon-7b 3.9 1.0 29.5 34.7 17.3 0.541
Llama-1-7b 10.7 3.0 26.4 36.1 19.0 0.510
Llama-2-7b 13.7 3.5 28.6 39.5 21.3 0.500
Llemma-7b 31.7 14.9 52.7 39.6 34.7 0.431
Mistral-7b 33.4 11.8 48.6 46.1 35.0 0.443
Qwen-7b 42.8 13.4 43.6 41.3 35.3 0.483
Yi-6b 31.1 6.5 33.6 40.8 28.0 0.483
@@
Llama-1-13b 13.6 4.1 28.6 39.0 21.3 0.487
Llama-2-13b 21.9 5.3 31.4 42.8 25.3 0.475
Qwen-14b 55.7 26.1 57.3 47.1 46.5 0.450
@@
Falcon-40b 23.0 4.0 31.8 43.0 25.5 0.482
Llama-1-30b 32.1 6.6 36.4 44.9 30.0 0.456
Llemma-34b 46.2 20.3 54.5 48.0 42.3 0.416
Yi-34b 61.4 16.6 52.3 50.8 45.3 0.421
@@
Deepseek-llm-67b 56.0 16.6 54.1 50.1 44.2 0.430
Jamba-52b 48.3 18.3 55.0 49.4 42.7 0.417
Llama-1-65b 42.5 10.0 41.8 45.3 34.9 0.441
Llama-2-70b 50.4 12.4 50.9 48.0 40.4 0.429
Mixtral-8x7b 53.4 23.7 64.1 54.4 48.9 0.394
Qwen-72b 72.6 36.2 71.4 52.3 58.1 0.415
@@ Appendix E Better Alignment between MMLU and Textbooks
The MMLU benchmark offers a comprehensive evaluation across various domains and tasks, leveraging datasets sourced from real-world examinations and textbooks. As a result, textbooks from various educational levels closely align with the benchmark’s distribution. We utilize a small collection of textbooks sourced from the internet as validation data. The results are shown in Figure  11. A stronger linear correlation is observed between the compression performance on the textbook data and the MMLU accuracy, characterized by a high Pearson correlation coefficient of -0.930.